OpenStudy (anonymous):

Find the first five terms of the sequence whose nth term is given by the following -4n-8

OpenStudy (gudden):

first term of this AP can be found out by using n = 1 in the given sequnce i.e. - 4 - 8 = - 12 and second term by keeping n = 2 -4(2) - 8 = -8 - 8 = -16 so to find the next 3 terms use n = 3,4 and 5